Problem 2 Two identical sound sources are placed on the x-axis, one at x=0, and other at x=40 m. They send identical, in-phasTwo identical sound sources are placed on the x-axis, one at x=0, and other at x=40 m. They send identical, in-phase sound wave of 3 m wave length out towards each other. Find the positions between, x=15 m and x=21 m where sound should be loudest.
Answer #1

\begin{aligned} & y_1=cos(\frac{2\pi}{\lambda}x)\\ & y_2=cos(\frac{2\pi}{\lambda}(40-x))\\ & y_1 + y_2 = max\\ &= cos(\frac{2\pi}{\lambda}x) +cos(\frac{2\pi}{\lambda}(40-x))\\ & =cos(\frac{2\pi}{\lambda}(40/2)).cos(\frac{2\pi}{\lambda}((40-2x)/2))\\ &= cos(\frac{2\pi}{3}(20)).cos(\frac{2\pi}{\lambda}(20-x)) \end{aligned}

= Maximum when x = 20m.